OpenStudy (anonymous):

what is the domain of the function f(x) = square root of x-3x^2

OpenStudy (anonymous):

x-3x^2 >= 0 || -3x^2 >= -x || 3 >= x/x^2 || 1/x >= 3 || 1 >= 3x || x >= 1/3